what volume of methanol should be used to make 1.25 L of a 21.0 % by volume solution
Answer #1

Answer:

Step 1: Explanation

Volume percentage is expressed as the ratio of  a volume of solute divided by the volume of the solution, multiplied by 100%

i.e   % v/v  = ( Volume of solute / Volume of solution ) x 100

Step 2: Extract the data from the question.

Given,

Percent (V/V) = 2%,

Volume of Solute = volume of methanol = we need to find

Volume of Solution = 1.25 L

Step 3:  Calculation of volume of methanol

By using above formula and sustituting the value

% v/v  = ( Volume of solute / Volume of solution ) x 100

=> 21 % = ( Volume of solute / 1.25 L ) x 100

=> Volume of solute =( 21 % x 1.25 L ) / 100 = 0.2625 L

Hence, the volume of methanol = 0.2625 L = 262.5 mL

[ because 1 L=1000 mL so, 0.2625 L = ( 0.2625 L  × ( 1000 mL / 1 L) = 262.5  mL ]